OpenStudy (anonymous):

Which of the following inequalities is graphed below? points (0,1) and (3,-1) I found the slope was -2/3 and it has to be with >,< with a equal to sign underneath it.

OpenStudy (anonymous):

the answers are \[A. y \ge -\frac{ 2 }{ 3 } x + 1 B. y \le -\frac{ 2 }{ 3 } x + 1\]